### v2

@@ 5. Unifying Bayesian, Frequentist, and Imprecise Reasoning
  • Fong and Holmes, (2021), who show that the negative posterior predictive density is a valid non-conformity score (i.e. it is an element of $\mathscr{F}$ in (4));
  • Lemma 9, which shows that map $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$ is jointly continuous on $Y^{n}\times Y=Y^{n+1}$ (and so also $\psi_{(y^{n},P)}(y^{n+1})$ is continuous, since we simply multiply $p_{P}(y_{n+1}\mid y^{n})$ by $-1$). This shows that $\psi_{(y^{n},P)}(y^{n+1})\equiv\psi_{P}(y^{n},y^{n+1})$ satisfies (ii).
Now, given our choice of non-conformity score, the Conformal Prediction Region, that is, the image $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})$ of the Conformal correspondence $\kappa_{\alpha}$, coincides with the $\alpha$-level set $\mathsf{QUANT}_{\alpha}(y^{n},P)$ of the posterior predictive distribution, for all $\alpha$ that satisfies (iii). To see this, notice that, by definition, a candidate $y\in Y$ lies in $\mathscr{R}^{\psi_{(y^{n},P)}}_{\alpha}(y^{n})\equiv\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})$ if and only if
|$|\{\,i:\psi_{(y^{n},P)}(y_{i})\geq\psi_{(y^{n},P)}(y_{n+1})\}|=|\{\,i:p_{P}(y_{i}\mid y^{n})\leq p_{P}(y_{n+1}\mid y^{n})\}|>\;k\coloneqq\lceil(n+1)\alpha\rceil.$
But among the $n$ scores $\psi_{(y^{n},P)}(y_{1}),\ldots,\psi_{(y^{n},P)}(y_{n})$, the event |$|\{i:\psi_{(y^{n},P)}(y_{i})\geq s\}|>k$ holds precisely when $s\leq\psi_{(y^{n},P)}(y_{(k)})$, where $\{\psi_{(y^{n},P)}(y_{(1)}),\ldots,\psi_{(y^{n},P)}(y_{(n)})\}$ are the ordered values of $\{\psi_{(y^{n},P)}(y_{1}),\ldots,\psi_{(y^{n},P)}(y_{n})\}$. Therefore,
$y\in\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})\quad\iff\quad\psi_{(y^{n},P)}(y_{n+1})\;\leq\;\psi_{(y^{n},P)}(y_{(k)}).$
Set $c(\alpha,y^{n},P)=\psi_{(y^{n},P)}(y_{(k)})$. Then,
$\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})=\{\,y_{n+1}\in Y:\psi_{(y^{n},P)}(y_{n+1})\leq c(\alpha,y^{n},P)\},$
that is, $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})$ is exactly the level set of the posterior predictive density at level $c(\alpha,y^{n},P)$.
By Caprio et al., 2025a (, Proposition 5), we also know that $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})=\mathsf{IHDR}_{\alpha}(\mathcal{M}(\overline{\Pi}))$, for all $\alpha$ that satisfies (iii). The remaining arrows follow directly from Theorem 5. ∎
@@
Theorem 11 is the central payoff of our categorical treatment, and perhaps the greatest joy (so far!) of looking at Conformal Prediction from a categorical angle. It confirms the intuition put forth in Martin, (2022) and Caprio et al., 2025a (, Section 5.2), that CP is a bridge between the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning. In the category UHCont, the model-based Bayesian and imprecise methods, and the model-free conformal construction, yield the same $\alpha$-level prediction region, thus showing a very profound connections between these three only apparently far apart prediction mechanisms.
@@
Notice also that Theorem 11 allows us to tie CP to the concept of e-posterior (Grünwald,, 2023), thus further testifying the central role of CP in modern Machine Learning.
@@
To see this, denote by $\Delta_{Y}^{\text{dens}}$ the space of finitely additive probability measures on $Y$ that admit a density with respect to a $\sigma$-finite dominating (atomless) measure $\lambda_{Y}$. Since we observe the realizations $y^{n}=(y_{1},\ldots,y_{n})$ of the first $n$ terms $\mathcal{Y}_{1},\ldots,\mathcal{Y}_{n}$ of an exchangeable process (as pointed out in Section 2.1), the likelihood density $p(y^{n}\mid\theta)\equiv\ell(y^{n}\mid\theta)$ can be written as (de Finetti,, 1974, 1975)
@@
Suppose that $\underline{p}$ is integrable. Then, the upper posterior density
$\overline{p}(\theta\mid y^{n})=\frac{\ell(y^{n}\mid\theta)\overline{p}(\theta)}{\underline{\ell}(y^{n})},$ (10)
where ¯$\underline{\ell}(y^{n})\coloneqq\int_{\Theta}\ell(y^{n}\mid\theta)\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)$, is an e-posterior à la Grünwald, (2023) if and only if $\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta)$, for all $\theta\in\Theta$.
We can interpret the necessary and sufficient condition in Proposition 12 as follows: The probability mass on which all priors agree is no larger than the minimum “peak plausibility” that any state enjoys under at least one prior.
@@
$\leq 1\iff\int_{Y^{n}}\underline{\ell}(y^{n})\lambda_{Y}^{\otimes n}(\text{d}y^{n})\leq\overline{p}(\theta),\quad\forall\theta\in\Theta.$
Now, notice that
@@
$\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta),\quad\forall\theta\in\Theta.$ (11)
This concludes the proof. ∎
@@
then the $\alpha$-level Imprecise Highest Density Region $\mathsf{IHDR}_{\alpha}(\mathcal{M}(\overline{\Pi}))$ in Theorem 11 is a prediction region that is associated with an e-posterior. The equality in (12) only holds if the upper posterior predictive density $\overline{p}(y_{n+1}\mid y^{n})$ collapses to a point mass.
@@
Assume (12). Then there exist points $y^{\ast}\in Y$ and $\theta^{\ast}\in\Theta$ such that, $\lambda_{Y}$- and $\lambda_{\Theta}$-almost surely,
$\overline{p}(y\mid y^{n})=\delta_{y^{\ast}}(y),\quad\overline{p}(\theta\mid y^{n})=\delta_{\theta^{\ast}}(\theta),\quad\ell(\,y\mid\theta^{\ast})=\delta_{y^{\ast}}(y).$
Proof.
Put
so (12) reads $\overline{\Pi}=\overline{\Pi}_{\mathrm{B}}$.
Equality of $\overline{\Pi}$ and $\overline{\Pi}_{\mathrm{B}}$ forces
Step 2 (A 00-$1$ law). If $\overline{\Pi}_{\mathrm{B}}(A),\overline{\Pi}_{\mathrm{B}}(B)\in(0,1)$, the right-hand side of (13) is strictly larger than the left, a contradiction. Hence
[…]
implying $\overline{\Pi}_{\mathrm{B}}=\delta_{y^{\ast}}$.
Step 4 (Collapse of the posterior and the likelihood). The mixture representation
$\overline{p}(y\mid y^{n})=\int_{\Theta}\ell(y\mid\theta)\,\overline{p}(\theta\mid y^{n})\,\lambda_{\Theta}(d\theta)=\delta_{y^{\ast}}(y)$
can hold only if $\overline{p}(\theta\mid y^{n})=\delta_{\theta^{\ast}}(\theta)$ for some $\theta^{\ast}$ and $\ell(\,\cdot\mid\theta^{\ast})=\delta_{y^{\ast}}$. ∎
Although Proposition 13 shows that the equality in (12) holds only under very restrictive conditions, it nevertheless hints at the possibility of an arrow in the diagram of Theorem 11 connecting $Y^{n}\times\Delta_{\Theta}^{\text{dens}}$ and $\mathscr{C}$ via an e-posterior-based method. [^11] To be precise, for such an arrow to be a morphism of $\mathbf{UHCont}$, we need to substitute $Y^{n}\times\Delta_{\Theta}^{\text{dens}}$ with $Y^{n}\times\mathscr{C}_{\Theta}^{\text{dens}}$, where $\mathscr{C}_{\Theta}^{\text{dens}}$ denotes the set of convex and weak-closed subsets of $\Delta_{\Theta}^{\text{dens}}$. In that case, $\mathsf{BCP}$ and $\mathsf{QUANT}_{\alpha}$ both take as input the pair $(y^{n},\{P\})$, where $\{P\}\subset\mathcal{P}_{\text{prior}}$. A deeper study of the broader interest of the e-posterior approach to Conformal Prediction within a category theory framework is left for future research.

### v3

@@ 4. Unifying Bayesian, Frequentist, and Imprecise Reasoning
First, we point out that Fong and Holmes, (2021) show that the negative posterior predictive density is a valid non-conformity score (i.e. it is an element of $\mathscr{F}$ in (4)). Fix then $\alpha\in(0,1)$ and write $r_{n}(y)\coloneqq p_{P}(y\mid y^{n})$. For $y\in Y$ define the full CP transducer
$\pi_{n}(y)=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}[\{r_{n,-i}(y_{i})\leq r_{n}(y)\}],$
with the convention $y_{n+1}=y$ and $r_{n,-(n+1)}(y_{n+1})\equiv r_{n}(y)$. Then $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})=\{y:\pi_{n}(y)\geq\beta_{n}(\alpha)\}$.
Step 1 (LOO reduction via localized bracketing). Set
[…]
Fix $\eta>0$ and define the $\alpha$-band
$\mathcal{B}_{\eta}(y^{n},P)\coloneqq\bigl{\{}y\in Y:|F(r_{n}(y);y^{n},P)-\alpha|\leq\eta\bigr{\}}.$
For every $y\in Y$,
$F_{n}\bigl{(}r_{n}(y)-\delta_{n};y^{n},P\bigr{)}\leq\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\bigl{[}r_{n,-i}(y_{i})\leq r_{n}(y)\bigr{]}\leq F_{n}\bigl{(}r_{n}(y)+\delta_{n};y^{n},P\bigr{)}.$
Hence, for $y\in\mathcal{B}_{\eta}(y^{n},P)$,
$\Bigl{|}\tfrac{1}{n}$ $\sum_{i=1}^{n}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}-F_{n}\bigl{(}r_{n}(y);y^{n},P\bigr{)}\Bigr{|}$
$\leq\sup_{\begin{subarray}{c}|t-r_{n}(y)|\leq\delta_{n}\end{subarray}}\bigl{|}F_{n}(t;y^{n},P)-F_{n}(r_{n}(y);y^{n},P)\bigr{|}$
$\leq\sup_{\begin{subarray}{c}|t-r_{n}(y)|\leq\delta_{n}\end{subarray}}\bigl{|}F(t;y^{n},P)-F(r_{n}(y);y^{n},P)\bigr{|}+2\sup_{u\in\mathbb{R}}\bigl{|}F_{n}(u;y^{n},P)-F(u;y^{n},P)\bigr{|}.$
By (ix), $c\mapsto F(c;y^{n},P)$ is continuous in a neighborhood of $c(\alpha,y^{n},P)$; Since $y\mapsto r_{n}(y)$ is continuous (Lemma 7) and $Y$ is compact, there is a deterministic modulus $\omega(\cdot)$ with
$\sup_{y\in\mathcal{B}_{\eta}(y^{n},P)}\sup_{|t-r_{n}(y)|\leq\delta}\bigl{|}F(t;y^{n},P)-F(r_{n}(y);y^{n},P)\bigr{|}\leq\omega(\delta),\qquad\omega(\delta)\downarrow 0\text{ as }\delta\downarrow 0.$
Together with (x).(b) we obtain the localized reduction
$\sup_{y\in\mathcal{B}_{\eta}(y^{n},P)}\Bigl{|}\tfrac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}-F\bigl{(}r_{n}(y);y^{n},P\bigr{)}\Bigr{|}=o_{\mathfrak{P}}(1).$
Finally, recall that
$\pi_{n}(y)=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}=\frac{n}{n+1}\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}+\frac{1}{n+1},$
so, since $n/(n+1)\to 1$,
$\sup_{y\in\mathcal{B}_{\eta}(y^{n},P)}\bigl{|}\pi_{n}(y)-F\bigl{(}r_{n}(y);y^{n},P\bigr{)}\bigr{|}=o_{\mathfrak{P}}(1).$
Step 2 (Decisions agree up to a shrinking $r$-tube). Let $c^{\star}\coloneqq c(\alpha,y^{n},P)$. By (ix), $c\mapsto F(c;y^{n},P)$ is strictly increasing near $c^{\star}$, so for any small $\delta>0$ define
$\gamma_{+}(\delta)\coloneqq F(c^{\star}+\delta;y^{n},P)-\alpha>0,\qquad\gamma_{-}(\delta)\coloneqq\alpha-F(c^{\star}-\delta;y^{n},P)>0,$
with $\gamma_{\pm}(\delta)\downarrow 0$ as $\delta\downarrow 0$. Recall that $\delta_{n}\coloneqq\sup_{y\in Y}\max_{i\leq n}|r_{n,-i}(y)-r_{n}(y)|\xrightarrow{\mathfrak{P}}0$. By LOO stability, for every $y\in Y$,
$F_{n}\bigl{(}r_{n}(y)-\delta_{n};y^{n},P\bigr{)}\leq G_{n}(y)\leq F_{n}\bigl{(}r_{n}(y)+\delta_{n};y^{n},P\bigr{)},$
where $G_{n}(y)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}$. Fix $\delta>0$ and write $c^{\star}=c(\alpha,y^{n},P)$. By (ix), $c\mapsto F(c;y^{n},P)$ is strictly increasing and continuous near $c^{\star}$. Thus, there is a modulus $\omega(\cdot)\downarrow 0$ such that for small $h$,
$\bigl{|}F(c^{\star}\pm\delta\pm h;y^{n},P)-F(c^{\star}\pm\delta;y^{n},P)\bigr{|}\leq\omega(|h|).$
Let $\mathscr{D}_{n}\coloneqq\sup_{t\in\mathbb{R}}\bigl{|}F_{n}(t;y^{n},P)-F(t;y^{n},P)\bigr{|}\xrightarrow{\mathfrak{P}}0$ by (x).(b). Then:
(Left inclusion). If $r_{n}(y)\geq c^{\star}+\delta$, then
$G_{n}(y)$ $\geq F_{n}\bigl{(}r_{n}(y)-\delta_{n}\bigr{)}\geq F\bigl{(}r_{n}(y)-\delta_{n}\bigr{)}-\mathscr{D}_{n}$
$\geq F\bigl{(}c^{\star}+\delta-\delta_{n}\bigr{)}-\mathscr{D}_{n}\geq\alpha+\gamma_{+}(\delta)-\omega(\delta_{n})-\mathscr{D}_{n},$
where $\gamma_{+}(\delta)\coloneqq F(c^{\star}+\delta;y^{n},P)-\alpha>0$. Hence,
$\pi_{n}(y)=\tfrac{n}{n+1}G_{n}(y)+\tfrac{1}{n+1}\geq\alpha+\gamma_{+}(\delta)-\omega(\delta_{n})-\mathscr{D}_{n}+\tfrac{1}{n+1}\geq\beta_{n}(\alpha),$
for all large $n$.
(Right inclusion). If $r_{n}(y)\leq c^{\star}-\delta$, then
$G_{n}(y)$ $\leq F_{n}\bigl{(}r_{n}(y)+\delta_{n}\bigr{)}\leq F\bigl{(}r_{n}(y)+\delta_{n}\bigr{)}+\mathscr{D}_{n}$
$\leq F\bigl{(}c^{\star}-\delta+\delta_{n}\bigr{)}+\mathscr{D}_{n}\leq\alpha-\gamma_{-}(\delta)+\omega(\delta_{n})+\mathscr{D}_{n},$
with $\gamma_{-}(\delta)\coloneqq\alpha-F(c^{\star}-\delta;y^{n},P)>0$, thus $\pi_{n}(y)<\beta_{n}(\alpha)$ for all large $n$.
Consequently, for any fixed $\delta>0$, with probability $\to 1$,
$\{y:r_{n}(y)\geq c^{\star}+\delta\}\subseteq\ \{y:\pi_{n}(y)\geq\beta_{n}(\alpha)\}\subseteq\ \{y:r_{n}(y)\geq c^{\star}-\delta\}.$ (11)
Thus any disagreement between $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})=\{y:\pi_{n}(y)\geq\beta_{n}(\alpha)\}$ and the target set $\{y:F(r_{n}(y);y^{n},P)\geq\alpha\}=\{y:r_{n}(y)\geq c^{\star}\}$ is confined to the “shrinking tube”
$\mathcal{T}_{\delta}(y^{n},P)\coloneqq\{y\in Y:|r_{n}(y)-c^{\star}|\leq\delta\}.$
Since $r_{n}$ is continuous on compact $Y$ (Lemma 7), the level sets $\{r_{n}\geq c^{\star}\pm\delta\}$ converge to $\{r_{n}\geq c^{\star}\}$ in Hausdorff distance as $\delta\downarrow 0$. Combined with (11), this yields the desired set convergence.
Step 3 (from inclusions to the target set). By (ix) and Lemma 7, the map $y\mapsto F(r_{n}(y);y^{n},P)$ is continuous on $Y$ at the decision level $\alpha$, and (see Lemma 8) the superlevel sets $\{y:F(r_{n}(y);y^{n},P)\geq t\}$ vary continuously in $t$ near $t=\alpha$. Therefore, as $\delta\downarrow 0$, the Hausdorff distance between $\{y:F(r_{n}(y);y^{n},P)\geq\alpha+\delta\}$ and $\{y:F(r_{n}(y);y^{n},P)\geq\alpha\}$ (and similarly for $\alpha-\delta$) tends to 0. Combining with the inclusions above yields
$d_{H}\left(\{y:\pi_{n}(y)\geq\beta_{n}(\alpha)\},\{y:F(r_{n}(y);y^{n},P)\geq\alpha\}\right)\xrightarrow{\mathfrak{P}}0,$
i.e. $d_{H}\bigl{(}\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)}),\mathsf{QUANT}_{\alpha}(y^{n},P)\bigr{)}\xrightarrow{\mathfrak{P}}0.$
Step 4 (Hausdorff convergence in probability). By Lemma 7 and (ix), the target $\alpha$-level set $\mathcal{Q}_{\alpha}(y^{n},P)\coloneqq\{y:F(r_{n}(y);y^{n},P)\geq\alpha\}=\mathsf{QUANT}_{\alpha}(y^{n},P)$ is a closed subset of the compact space $Y$. The sandwich above implies that, on events with probabilities $\to 1$, the inner and outer deviations of $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})$ from $\mathcal{Q}_{\alpha}(y^{n},P)$ shrink to 0. Therefore,
$d_{H}\left(\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)}),\mathcal{Q}_{\alpha}(y^{n},P)\right)\xrightarrow{\mathfrak{P}}0.$
Step 5 (Top route). Under consonance, Theorem 5 gives $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$ for any score. Composing with $\mathsf{BCP}$ yields the second convergence in probability. ∎
@@
Theorem 9 is the central payoff of our categorical treatment, and perhaps the greatest joy (so far!) of looking at Conformal Prediction from a categorical angle. It confirms the intuition put forth in Martin, (2022) and Caprio et al., 2025a (, Section 5.2), that CP is a bridge between the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning. When seen as u.h.c. correspondences, the model-based Bayesian and imprecise methods, and the model-free conformal construction, yield (asymptotically) the same $\alpha$-level prediction region, thus showing a very profound connections between these three only apparently far apart prediction mechanisms.
4.1. Conformal Prediction and e-Posteriors
@@
Theorem 9 allows us to tie CP to the concept of e-posterior (Grünwald,, 2023) as well, thus further testifying the central role of CP in modern Machine Learning.
@@
To see this, recall that $\Delta_{Y}^{\text{dens}}$ denotes the space of finitely additive probability measures on $Y$ that admit a density with respect to a $\sigma$-finite dominating (atomless) measure $\lambda_{Y}$. Since we observe the realizations $y^{n}=(y_{1},\ldots,y_{n})$ of the first $n$ terms $\mathcal{Y}_{1},\ldots,\mathcal{Y}_{n}$ of an exchangeable process (as pointed out in Section 2.1), the likelihood density $p(y^{n}\mid\theta)\equiv\ell(y^{n}\mid\theta)$ can be written as (de Finetti,, 1974, 1975)
@@
Suppose that $\underline{p}\in L^{1}(\lambda_{\Theta})$, that $\underline{\ell}(y^{n})\coloneqq\int_{\Theta}\ell(y^{n}\mid\theta)\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)>0$, for all $y^{n}\in Y^{n}$, and that the upper prior density $\overline{p}(\theta)\in(0,\infty)$, for all $\theta\in\Theta$. Then the upper posterior
$\overline{p}(\theta\mid y^{n})=\frac{\ell(y^{n}\mid\theta)\overline{p}(\theta)}{\underline{\ell}(y^{n})}$ (12)
is an e-posterior in the sense of Grünwald, (2023) if and only if
$\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta),\quad\text{for all }\theta\in\Theta.$ (13)
We can interpret the necessary and sufficient condition in Proposition 10 as follows: The probability mass on which all priors agree is no larger than the minimum “peak plausibility” that any state enjoys under at least one prior. Since Proposition 10 is of general significance (the IP and e-values communities may be interested in it), we leave its proof in the main portion of our paper.
@@
$\leq 1\iff\int_{Y^{n}}\underline{\ell}(y^{n})\lambda_{Y}^{\otimes n}(\text{d}y^{n})\leq\overline{p}(\theta),$
for all $\theta\in\Theta$. Now, notice that
@@
$\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta),$ (14)
for all $\theta\in\Theta$. This concludes the proof. ∎
@@
then the $\alpha$-level Imprecise Highest Density Region $\mathsf{IHDR}_{\alpha}(\mathcal{M}(\overline{\Pi}))$ in Theorem 9 is a prediction region that is associated with an e-posterior. The equality in (15) only holds if the upper posterior predictive density $\overline{p}(y_{n+1}\mid y^{n})$ collapses to a point mass. We will denote
so that (15) reads $\overline{\Pi}=\overline{\Pi}_{\mathrm{B}}$. The following result is proved in Appendix B.
@@
Assume (15), and relax the assumption that $\lambda_{Y}$ is atomless. Then, there exists $y^{\ast}\in Y$ such that
$\overline{\Pi}_{\mathrm{B}}(\cdot)=\delta_{y^{\ast}}(\cdot)\quad\text{and}\quad\overline{P}\left(\{\theta:L_{\theta}=\delta_{y^{\ast}}\}\middle|y^{n}\right)=1.$
In particular, if $\overline{P}(\cdot\mid y^{n})\ll\lambda_{\Theta}$, then necessarily $\lambda_{Y}(\{y^{\ast}\})>0$ and there exists a density $\overline{p}(\cdot\mid y^{n})\in L^{1}(\lambda_{\Theta})$ supported on $\Theta^{\ast}\coloneqq\{\theta:L_{\theta}=\delta_{y^{\ast}}\}$, i.e.
$\overline{p}(\theta\mid y^{n})=0\quad\text{for }\lambda_{\Theta}\text{-almost all }\theta\notin\Theta^{\ast},\qquad\int_{\Theta^{\ast}}\overline{p}(\theta\mid y^{n})\lambda_{\Theta}(d\theta)=1.$
On $Y$ we have
$\overline{p}(y\mid y^{n})=\frac{\mathbbm{1}(y=y^{\ast})}{\lambda_{Y}(\{y^{\ast}\})}.$
If, in addition, the model is identifiable (the map $\theta\mapsto L_{\theta}$ is injective), then $\overline{P}(\cdot\mid y^{n})=\delta_{\theta^{\ast}}$.
If $\lambda_{Y}$ is atomless and $\overline{p}_{P}(\cdot\mid y^{n})\in L^{1}(\lambda_{Y})$, then (15) cannot hold for all $A\in\Sigma_{Y}$. Indeed, (15) would force $\overline{\Pi}_{\mathrm{B}}=\delta_{y^{\ast}}$, which is not absolutely continuous with respect to an atomless $\lambda_{Y}$. Although Proposition 11 shows that the equality in (15) holds only under very restrictive conditions, it nevertheless hints at the possibility of an arrow in the diagram of Theorem 9 connecting $Y^{n}\times\Delta_{\Theta}^{\text{dens}}$ and $\mathscr{C}$ via an e-posterior-based method. [^13] To be precise, for such an arrow to be a morphism of $\mathbf{UHCont}$, we need to substitute $Y^{n}\times\Delta_{\Theta}^{\text{dens}}$ with $Y^{n}\times\mathscr{C}_{\Theta}^{\text{dens}}$, where $\mathscr{C}_{\Theta}^{\text{dens}}$ denotes the set of convex and weak-closed subsets of $\Delta_{\Theta}^{\text{dens}}$. In that case, $\mathsf{BCP}$ and $\mathsf{QUANT}_{\alpha}$ both take as input the pair $(y^{n},\{P\})$, where $\{P\}\subset\mathcal{P}_{\text{prior}}$. A deeper study of the broader interest of the e-posterior approach to Conformal Prediction within a category theory framework is left for future research.